two integers are such that the greater integer added to twice the smaller integer gives 52. find the greater integer.pls show the how to answer such questions
Answer by mananth(16949)

two integers are such that the greater integer added to twice the smaller integer gives 52. find the greater integer.pls show the how to answer such questions
..
let integers be x, x+1
..
greater integer added to twice smaller : 2x+x+1=52
3x+1=52
add-1
3x+1-1=52-1
3x=51
/3
x= 17
..
17 , 18
...
m.ananth@hotmail.ca